Shares in Tichibo are selling for $55 per share. There are one million shares outstanding. What will be the share price in each of the following situations? Ignore taxes. (Round your answers to the nearest cent.)

a. The stock splits five-for-four.
  Price $

       

b. The company pays a 24% stock dividend.
  Price $

                

c. The company repurchases 127,000 shares.
  Price $
Answer #1

a. 5 for 4 stock split means that 5 shares will be issued for every 4 held.

Price per share = 55*4/5 = $44

B. 24% stock dividend means that 0.24 shares will be issued as dividend for every share held.

Price per share = 55/1.24 = $44.35

C. Since the shares have been repurchased at market price, no change in the price per share

Price per share = $55